Suppose you've just purchased a new warehouse for $4,500,000. To finance the purchase, you've arranged for a 30-year mortgage for 80% of the $4,500,000 purchase price. The monthly payment on this loan will be $27,500. What is the APR on this loan? What is the EAR?
Answer #1

Solution:-

Loan amount = 45, 00,000 * 80% = $36, 00,000

PV = 36, 00,000

N = 30 * 12 = 360

PMT = -27,500

FV = 0

Use excel function RATE to find APR=RATE (nper, pmt, pv, fv)

Monthly rate=RATE (360,-27500, 36, 00, 00 ,0)=0.70%

APR:-

APR = 12 * (0.70%)

         = 8.4%

EAR:-

EAR = (1 + Interest Rate) ^12 – 1

         = (1 + 0.007) ^12 – 1

         = 0.0873

         = 8.73